OpenStudy (anonymous):

Which mixed number is equal to 7.6? A. B. C. D.

OpenStudy (anonymous):

Well, we know our whole number is 7 , so we have \(7\dfrac{?}{?}\)

OpenStudy (anonymous):

Now we are left with 0.6. Do you know what 0.6 is in fraction form>

OpenStudy (anonymous):

no

ganeshie8 (ganeshie8):

multiply by \(\dfrac{10}{10}\)

ganeshie8 (ganeshie8):

\[\large 0.6 = 0.6\times \dfrac{10}{10} = \dfrac{6}{10}\] cancel any common factors

OpenStudy (anonymous):

so what is the final answer

ganeshie8 (ganeshie8):

what do you get after simplifying \(\large \dfrac{6}{10}\) ?

OpenStudy (anonymous):

1

ganeshie8 (ganeshie8):

\[\dfrac{6}{10} = \dfrac{2\times 3}{2\times 5}= \dfrac{3}{5}\]

OpenStudy (anonymous):

it has to be a mixed number

ganeshie8 (ganeshie8):

\[7.6 = 7 + 0.6 = 7 + \dfrac{3}{5}\] for mixed number just remove the + sign \[7\frac{3}{5}\]

OpenStudy (anonymous):

that not choice